OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4).

OpenStudy (anonymous):

y =4/3 x + 8 y =4/3 x + 3 y = −3/4x + 8 y = −3/4x + 3

OpenStudy (anonymous):

this is basically the same type of problem as the last. again we need to find the slope of that given line. I already started for you to find slope. solve for y in that last line...